OpenStudy (anonymous):

Is it possible to factor f(x) = x^2 - 4

OpenStudy (anonymous):

x^2-4 = x^2-2^2 = (x+2)(x-2)

OpenStudy (anonymous):

So, what do you think now. is it possible?

OpenStudy (anonymous):

yes i think so... man I am really bad at math :(

OpenStudy (anonymous):

thank you sooooooo much

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

s if you have an equation x^4 - 4 it would factor to (x^2+2) (x^2-2)